OpenStudy (anonymous):

the function defined by f(x)=x^3-3x^2 for all real numbers x has a relative maximum at x=

OpenStudy (anonymous):

x^3-3x^2=x^2(x-3) one can use derivative test and check for x=-sqrt(2)

hartnn (hartnn):

to get relative max., differentiate f(x) and equate it to 0 that is, f'(x)=0 this function has relative maximum at x=0.

OpenStudy (anonymous):

as well as x=+sqrt(2) and x=0

hartnn (hartnn):

not at x=sqrt 2... only at x=0